OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form.

pooja195 (pooja195):

use the point slope formula :) Y-Y1=M(X-X1) then just plug in the numbers =:) (x1,y1) m=slope

OpenStudy (anonymous):

y-4 = -1* (x-2) y-4=-x +2 x+y=6